OpenStudy (anonymous):

Can sumone help me with this polynomial problem: Which of the following is the LCM of the polynomials? (x^2 + 10x + 9) and (x^2 - 1) A. (x + 1) B. (x + 1)( x + 9) C. (x + 1)(x - 1) D. (x + 1)(x - 1)(x + 9)

OpenStudy (anonymous):

x^2+x+9x+9 = x(x+1)+9(x+1) = (x+9)(x+1) x^2-1 = (x+1)(x-1) hence the answer is a) (x+1)

OpenStudy (anonymous):

oh sorry the answer is D) (x+1)(x-1)(x+9) i have to find lcm forgot :0

OpenStudy (anonymous):

my answer is correct

OpenStudy (anonymous):

LCM of (x^2 + 10x +9) = (x+9)(x+1) LCM of (x^2 - 1) = (x + 1)(x - !) Hence, the LCM of both of them is (x + 1) (x - 1) (x + 9)

OpenStudy (anonymous):

okay thnk u guys very much.

OpenStudy (anonymous):

You are most welcome dear

OpenStudy (anonymous):

I like u rprofile pic lol

OpenStudy (anonymous):

Thank You